### v1

# Piecewise Polynomial Tensor Network Quantum Feature Encoding

Abstract
This work introduces a novel method for embedding continuous variables into quantum circuits via piecewise polynomial features, utilizing low-rank tensor networks. Our approach, termed Piecewise Polynomial Tensor Network Quantum Feature Encoding (PPTNQFE), aims to broaden the applicability of quantum algorithms by incorporating spatially localized representations suited for numerical applications like solving partial differential equations and function regression. We demonstrate the potential of PPTNQFE through efficient point evaluations of solutions of discretized differential equations and in modeling functions with localized features such as jump discontinuities. Although promising, challenges such as unexplored noise impact and design of trainable circuits remain. This study opens new avenues for enhancing quantum models with novel feature embeddings and leveraging TN representations for a wider array of function types in quantum machine learning.
Keywords Quantum Machine Learning  $\cdot$ Feature Encoding  $\cdot$ Tensor Networks  $\cdot$ Piecewise Polynomial  $\cdot$ Splines  $\cdot$ Differential Equations  $\cdot$ Finite Element Methods
1 Introduction
In the quest for viable quantum computing, Quantum Machine Learning (QML) has emerged as a rapidly expanding domain of research. The aspiration is that a model could discern patterns more effectively by navigating the quantum state space, as opposed to relying solely on classical states. It appears reasonable to assert that this methodology is advantageous for data emanating from inherently quantum processes, and there is now substantial evidence supporting this claim Huang et al. (2021a, 2022). Nonetheless, given that the majority of data-driven applications do not operate within the quantum realm, the pertinent question remains: does QML hold utility for classical data and problems?
1.1 Quantum Advantage in Quantum Machine Learning
A central theme in Quantum Machine Learning (QML), and in quantum computing at large, is the pursuit of the elusive “quantum advantage” (QA). This term is broadly defined as the superior performance of a quantum model over its best classical counterpart. Despite the numerous claims of achieving QA, practical applications remain conspicuously absent Lee et al. (2023); Abbas et al. (2023); Montanaro and Pallister (2016); Ali and Kabel (2023).
Theoretically, there exist rigorous demonstrations of QA in specific learning contexts Liu et al. (2021); Muser et al. (2023); Yamasaki et al. (2023). However, these proofs are often applicable to contrived problems lacking real-world relevance. On the practical front, the capabilities of contemporary quantum devices are limited by their small scale and significant noise levels, rendering them inadequate for large-scale demonstrations. Efficiently scaling quantum model training remains debatable Abbas et al. (2023), yet our focus here is more on generalization, especially through the lens of feature embeddings.
The reevaluation of QA as the primary goal for QML has been insightfully suggested by Schuld and Killoran (2022). A more constructive approach may involve exploring the unique characteristics of quantum models and identifying effective foundational elements. Despite this, there remains a significant focus within the literature on developing intricate hybrid models. However, even among these advanced models, the underlying building block often remains the relatively straightforward concept of rotation-based parametrized quantum circuits (PQCs), see Figure 2.
The rationale for employing PQCs often hinges on the large size of the quantum state space. However, these circuits often produce models similar to Fourier series, which may not fully take advantage of the exponentially large Hilbert space and can be emulated classically Schuld et al. (2021); Kordzanganeh et al. (2023); Shin et al. (2023), or they may be overly expressive, hindering effective training and generalization Kübler* et al. (2021); Huang et al. (2021b); Jerbi et al. (2023); Cerezo et al. (2023). Furthermore, the impact of statistical and hardware-induced errors is frequently overlooked, potentially undermining any claims of enhanced expressiveness or precision.
1.2 Inductive Bias and Feature Encoding
The efficacy of a model extends beyond mere expressiveness; it depends on the synergy between sufficient expressiveness and the appropriate, problem-specific inductive bias. Recent innovative studies have explored the potential to imbue models with biases that are distinctively quantum in nature Bowles et al. (2023); Gili et al. (2023); Nguyen et al. (2022); Schatzki et al. (2024).
A pivotal aspect common to both classical and quantum models is feature encoding, the process by which input features are integrated into the model’s latent spaces, enabling the training algorithm to detect patterns. Several promising methods have emerged to embed biases through feature encoding in QML Peters and Schuld (2023); Cong et al. (2019); Meyer et al. (2023); Umeano et al. (2023, 2024). Yet, the prevalent practice in QML models, especially those dealing with classical data, remains the use of rotation-based encodings. This approach inherently predisposes the model to identify wave-like, global, periodic patterns, which, while beneficial for certain tasks, constitutes a constraint for many others.
1.3 This Work
To broaden the spectrum of quantum feature embeddings (QFEs), this work introduces an approach to efficiently integrate piecewise polynomial features into quantum models. Piecewise polynomial approximations are widely utilized in various numerical applications, including interpolation and solving partial differential equations (PDEs), as well as in classical deep learning frameworks like networks employing rectified linear units (ReLUs). Spatially localized representations, such as those provided by piecewise polynomials, are often more suitable for certain applications compared to global trigonometric polynomials Schuld et al. (2021) or Chebyshev polynomials Kyriienko et al. (2021); Williams et al. (2023).
Our approach leverages an intermediate tensor network (TN) representation, which makes it adaptable to any function that can be represented in low-rank format Oseledets (2013); Ali and Nouy (2023). We denote these embeddings as Piecewise Polynomial Tensor Network Quantum Feature Encodings (PPTNQFE). It is crucial to note that these embeddings are not inherently superior or inferior to others and do not confer any QA. Through toy examples, we elucidate the scenarios where PPTNQFEs are advantageous and when they are not.
The structure of this paper is as follows. Section 2 reviews piecewise polynomials and the properties required for embedding. Section 3 discusses TNs and functional TN representations. Section 4 details the methodology for expressing piecewise polynomial bases as low-rank TNs and their subsequent translation into quantum circuits. Section 5 presents two important applications: point evaluations of PDE solutions and function learning (regression). The paper concludes in Section 6. The implementation and examples provided are confined to one spatial dimension, with further details in Section 7. While extending these methods efficiently to higher dimensions is theoretically feasible, it would entail a more involved implementation.
2 Piecewise Polynomials
Within a bounded domain $\Omega\subset\mathbb{R}^{d}$ and a partition $\mathcal{T}_{\tilde{N}}:=\{T_{0},\ldots,T_{\tilde{N}-1}\}$ that collectively covers $\overline{\Omega}=\bigcup_{i}T_{i}$, functions $S_{N}:=\{\varphi_{0},\ldots,\varphi_{N-1}\}$ are said to be piecewise polynomial if the restriction $\varphi_{k}|_{T_{i}}$ is a polynomial for every $k$ and $i$. This framework is typical in finite element methods (FEM) Braess (2007). Similarly, in learning, (leaky) ReLU networks yield piecewise polynomial functions Yarotsky (2017). For numerical stability and network trainability, these functions are usually required to satisfy global regularity, $\varphi_{k}\in C^{m}(\Omega)$, with $m\geq 0$. This in particular excludes discontinuous TN polynomial features as in Ali and Nouy (2023). While TN approximations might offer implicit regularization through low ranks (see examples in Nouy et al. (2023)), embedding discontinuous features into quantum circuits complicates training.
Two essential properties are stipulated for this set of functions:
  1. (I)
    For any given partition $\mathcal{T}_{\tilde{N}}$, each element $T_{i}$ is predetermined, i.e., parameter-independent, and its coordinates can be computed in $\mathcal{O}(1)$. This excludes replicating the ReLU approximations as in Yarotsky (2017).
  2. (II)
    The functions in $S_{N}$ have local support:
    #$\#\left\{k=0,\ldots,N-1:|\operatorname{supp}\varphi_{k}\cap\operatorname{supp}\varphi_{i}|>0\right\}=\mathcal{O}(1),$
    for any $i$. In other words, for any given $x\in\overline{\Omega}$, there is only a constant number of $\varphi_{k}$’s such that $\varphi_{k}(x)\neq 0$.
The locality property (II) leads to sparse embedding vectors. Instead of sparsity, we could require any set of functions that are representable with low-rank TNs such as, e.g., “saw-tooth” functions, see Ali and Nouy (2023).
Additionally, though not mandatory, it is beneficial for $S_{N}$ to form a partition of unity:
$\sum_{k}\varphi_{k}(x)=1,\quad\forall x\in\overline{\Omega}.$ (III)
In this work, we will explicitly derive and implement an embedding for a 1D basis on the interval $\overline{\Omega}=[-1,1]$. The basis consists of piecewise linear hat functions
$\hat{\varphi}_{k}(x)=\begin{cases}\frac{x-x_{k-1}}{h},&\quad\text{for }x_{k-1}\leq x<x_{k},\\ \frac{x_{k+1}-x}{h},&\quad\text{for }x_{k}\leq x<x_{k+1},\\ 0,&\quad\text{otherwise},\end{cases}$
for uniformly spaced $x_{k}$, where $h:=1/(N-1)$. The placing of the nodes $x_{k}$ will depend on the application:
  • In Section 5.1, $h=1/(2^{n}-1)$, $x_{0}=-1+h$, $x_{2^{n}-1}=1-h$, where $n$ is the number of qubits.
  • In Section 5.2, $x_{0}=-1$, $x_{2^{n}-1}=1$.
A plot of the hat functions is presented in Figure 1. These satisfy conditions (I), (II) and (III).
Figure 1: 1D piecewise linear hat basis for a grid of 8 uniformly spaced points, including the boundary $x\in\{\pm 1\}$
3 Tensor Networks
Tensor networks are an implicit, parsimonious representation of high-order tensors. They are most commonly known for their effectiveness in modeling high-dimensional quantum systems with low entanglement. For details on TN formats and diagrams, see Biamonte and Bergholm (2017). In this work, we will be using the matrix product state (MPS) format Schollwöck (2011), also known under the name tensor train (TT) format Oseledets (2011), see Figure 2.
Figure 2: On the top left, an example of a rotation-based exponential encoding. These embeddings can be repeated, or applied to multidimensional features $x$ with multi-qubit rotations depending on products of features $R_{ZZ}(x_{i}x_{j})$, see also Havlíček et al. (2019). On the top right, an encoding of the features into a matrix product state (MPS) with maximum rank 2 and converting the latter to a quantum circuit. On the bottom, an example of a variational MPS layer that we will require in Section 5.2.
A crucial feature of TNs is that they can be converted into quantum circuits. This was done for MPS in Ran (2020). For this work, the implementation was based on the details specified in Lubasch et al. (2020), more details can be found in Section 7. The width of the multi-qubit gates depends logarithmically on the maximum MPS rank. For the MPS embedding presented in this work, the maximum rank is 2 and thus we only require 2-qubit gates for the conversion in Figure 2.
For loading sparse vectors onto a quantum computer, we do not need TNs and could, in principle, use other methods as in Gleinig and Hoefler (2021). However, TNs offer advantages for function representations and approximations that could potentially allow the method to be generalized to other sets of functions. There is an extensive theory behind TN properties for function approximation (see, e.g., Oseledets (2013); Ali and Nouy (2023); Kazeev et al. (2017)), numerical methods (see, e.g., Kazeev et al. (2017); Bachmayr and Kazeev (2020); Haberstich et al. (2023); Nouy et al. (2023); Ritter et al. (2024)), and applications to quantum computing (see, e.g., Holmes and Matsuura (2020); Melnikov et al. (2023); Gourianov et al. (2022); Miyamoto and Ueda (2023); Iaconis et al. (2024)). Moreover, as mentioned in Section 2, we could also load nonlocal function sets such as saw-tooth functions.
4 Methodology
Prior to transferring the basis functions from Section 2 to a quantum computer, a transformation $\varphi_{k}=g\circ\hat{\varphi}_{k}$ might be necessary. For instance, in Section 5.2, we apply the transformation $\varphi_{k}=\sqrt{\hat{\varphi}_{k}}$. We only require this transformation to preserve the locality property (II). Additionally, normalization might require monitoring the norm
$n_{\Phi}(x):=\left(\sum_{k}\varphi_{k}^{2}(x)\right)^{1/2}.$
Given the property (II), $n_{\Phi}(x)$ can be computed in $\mathcal{O}(1)$ for any $x\in\overline{\Omega}$. For the case $\varphi_{k}=\sqrt{\hat{\varphi}_{k}}$, the property (III) ensures $n_{\Phi}(x)=1$, facilitating automatic normalization of the basis vector.
We have the basis vector
$\Phi(x):=\begin{bmatrix}\varphi_{0}(x)&\\ \vdots&\\ \varphi_{2^{n}-1}(x)&\end{bmatrix},$
and we want to prepare the corresponding normalized quantum state,
|$\ket{\Phi(x)}:=\frac{1}{n_{\Phi}(x)}\sum_{k=0}^{2^{n}-1}\varphi_{k}(x)\ket{k},$
where |$\ket{k}$ are the computational basis states. The basis vector $\Phi(x)$ is initially represented as an MPS. For $n$ qubits, we write an integer index in binary form $k=i_{0}i_{1}\cdots i_{n-1}$, where we use the Big Endian convention (the most significant bit is $i_{0}$). Then, we seek a representation
$\Phi(x)=\left(\varphi_{i_{0}\cdots i_{n-1}}(x)\right)_{i_{0}=0,\ldots,i_{n-1}=0}^{1}=C^{0}(x)[\cdot,i_{0},\cdot]\cdots C^{n-1}(x)[\cdot,i_{n-1},\cdot],$
where the order-3 cores have dimensions $C^{j}(x)\in\mathbb{C}^{r_{j}\times 2\times r_{j+1}}$ and open boundary conditions $r_{0}=r_{n}=1$.
In order for the complexity of the embedding circuits to remain small, it is important the ranks $r_{j}$ remain small as well. There are two properties crucial for this: the number of non-zero entries in $\Phi(x)$ and their relative position with respect to each other, i.e., the choice of ordering for the basis functions with respect to $k=i_{0}\cdots i_{n-1}$. Alternatively, as mentioned in Section 2, we could use non-local bases that have a low-rank representation. We demonstrate how to represent the 1D hat basis from Section 2 as a TN.
For any $x\in[-1,1]$, there are at most two adjacent non-zero values,
$\Phi(x):=\begin{bmatrix}0&\\ \vdots&\\ \varphi_{k}(x)=:a&\\ \varphi_{k+1}(x)=:b&\\ \vdots&\\ 0&\end{bmatrix},$
The MPS representation varies depending on whether the index $k$ is even or odd, with a detailed explanation provided for the latter. Let
$k$ $=i_{0}\cdots i_{j^{*}-1}0_{j^{*}}1_{j^{*}+1}\cdots 1_{n-2}1_{n-1},$
$k+1$ $=i_{0}\cdots i_{j^{*}-1}1_{j^{*}}0_{j^{*}+1}\cdots 0_{n-2}0_{n-1},$
where $j^{*}$ is the position of the least significant 00-bit in $k$. The nonzero entries of the MPS cores are given as
$C^{j}(x)$ $\in\mathbb{R}^{1\times 2\times 1},$
$C^{j}(x)[0,i_{j},0]$ $=1,$
for $j=0,\ldots,j^{*}-1$,
$C^{j}(x)$ $\in\mathbb{R}^{1\times 2\times 2},$
$C^{j}(x)[0,0,1]$ $=1,$
$C^{j}(x)[0,1,0]$ $=1,$
for $j=j^{*}$,
$C^{j}(x)$ $\in\mathbb{R}^{2\times 2\times 2},$
$C^{j}(x)[0,0,0]$ $=1,$
$C^{j}(x)[1,1,1]$ $=1,$
for $j=j^{*}+1,\ldots,n-2$, and
$C^{n-1}(x)$ $\in\mathbb{R}^{2\times 2\times 1},$
$C^{n-1}(x)[0,0,0]$ $=b,$
$C^{n-1}(x)[1,1,0]$ $=a.$
The case of even $k$ is similar. Since the maximum rank is $2$, converting to a quantum circuit as in Figure 2. we require only $2$-qubit gates.
Remark 4.1
We stress the difference from the work in Holmes and Matsuura (2020); Melnikov et al. (2023); Miyamoto and Ueda (2023); Iaconis et al. (2024): here we load the set of functions $\{\varphi_{k}(x)\}_{k=0}^{N-1}$, evaluated at a point $x\in\overline{\Omega}$, instead of loading a discrete representation of a single function $\{f_{k}=f(x_{k})\}_{k=0}^{N-1}$.
5 Applications
In this section, we dive into two illustrative applications of the PPTNQFE method that we have introduced. The first application focuses on the efficient point evaluation of solutions to differential equations. In this scenario, the discretized and normalized solution is encoded in a quantum state, with the basis coefficients represented within the quantum amplitudes. The second application we explore is regression, specifically the supervised learning of functions depending on continuous variables from given data samples. Here, PPTNQFE allows, e.g., the approximation of jump functions in a particularly simple manner.
5.1 Differential Equations
The prevalent methodologies for solving PDEs on quantum computers involve mapping the problem’s degrees of freedom into the amplitudes of a quantum state. This strategy has been applied to various types of PDEs, including linear and nonlinear, as well as stationary and time-dependent cases. In both quantum and classical digital computing, the initial step involves discretizing the infinite-dimensional PDE problem into a finite-dimensional system. Within the framework of FEM, a widely used discretization technique, the approximate solution is represented as
$u\approx\sum_{k}\boldsymbol{u}_{k}\varphi_{k},$
where $\varphi_{k}$ are piecewise polynomials as described in Section 2, and $\boldsymbol{u}=(\boldsymbol{u}_{k})_{k}$ is the solution vector containing the degrees of freedom.
The next step involves solving the potentially nonlinear system
$\boldsymbol{A}(\boldsymbol{u})=\boldsymbol{f},$ (1)
where $\boldsymbol{A}$ and $\boldsymbol{f}$ represent the differential operator, the right-hand side (RHS), and possibly the boundary conditions, all formulated in the basis $S_{N}$. On a quantum computer, solving (1) can be approached using algorithms such as the HHL algorithm Harrow et al. (2009); Childs et al. (2017) or variational methods Bravo-Prieto et al. (2023); Lubasch et al. (2020). The outcome is a quantum state
|$\ket{\boldsymbol{u}}:=\frac{\boldsymbol{u}}{\|\boldsymbol{u}\|}=\frac{1}{\|\boldsymbol{u}\|}\sum_{k}\boldsymbol{u}_{k}\ket{k},$
or an approximation thereof.
In quantum computing, unlike classical computing, the coefficients of $\boldsymbol{u}$ are encoded in the state’s amplitudes and are not directly accessible. Typically, one can only measure observables related to |$\ket{\boldsymbol{u}}$. However, for point evaluations of the solution $u(x)$, one can use the relationship
$u(x)=\sum_{k}\boldsymbol{u}_{k}\varphi_{k}(x)=\|\boldsymbol{u}\|n_{\Phi}(x)\Re\braket{\boldsymbol{u}}{\Phi(x)}.$
where $n_{\Phi}(x)$ is either $1$ or can be efficiently calculated in $\mathcal{O}(1)$. Utilizing preparation procedures for |$\ket{\Phi(x)}$ and |$\ket{\boldsymbol{u}}$, and employing techniques like the Hadamard test, one can efficiently estimate $\Re\braket{\boldsymbol{u}}{\Phi(x)}$ and thus $u(x)$.
An illustrative example is provided by the Dirichlet problem for a second-order differential equation
$-u^{\prime\prime}(x)$ $=f(x),$ $\quad x\in(-1,1),$
$u(x)$ $=0,$ $\quad x\in\{\pm 1\}.$
The associated linear system, represented in matrix form,
$\boldsymbol{A}\boldsymbol{u}=\frac{1}{h}\begin{pmatrix}2&-1&0&\dots&\\ -1&2&-1&\dots&\\ &&\ddots&&\\ &&&&-1\\ &&&-1&2\end{pmatrix}\boldsymbol{u}=\boldsymbol{f}.$
can be solved using a variational quantum algorithm. We use |$\ket{\boldsymbol{f}}=H^{\otimes n-1}\otimes X\ket{0}$ for the RHS, which corresponds to a discontinuous jump function. Details of all required circuits can be found in Ali and Kabel (2023). The resulting quantum state |$\ket{\boldsymbol{u}}$ approximates the normalized solution, and point evaluations of $u(x)$ can be efficiently computed and compared with the exact solution. Figure 3 plots the exact normalized solution, the quantum approximation |$\ket{\boldsymbol{u}}$ and the point evaluations $\frac{1}{\|\boldsymbol{u}\|}u(x)$.
Figure 3: We plot the exact classical solution of the linear system of equations, the quantum state amplitudes on 5 qubits prepared with a variational solver as in Ali and Kabel (2023) and the quantum solution evaluated at 10 points.
5.2 Regression
In the context of supervised learning, we are presented with a training set $\mathcal{S}_{M}:=\left\{(x_{0},y_{0}),\ldots,(x_{M},y_{M})\right\}\in(X\times Y)^{M}$ consisting of independently and identically distributed samples from an unknown distribution $\mathcal{P}(X\times Y)$ and a hypothesis class $\mathcal{H}$. The objective is to identify a model $h\in\mathcal{H}$, $h:X\rightarrow Y$, that minimizes the empirical risk
$R_{M}(h):=\frac{1}{M}\sum_{i}L(h(x_{i}),y_{i}),$
where $L:Y\times Y\rightarrow\mathbb{R}$ is a specified loss function.
In QML, hypothesis classes often derive from PQCs:
|$\ket{\psi(x,\theta)}$ $:=U(x,\theta)\ket{0},$
$h(x,\theta)$ $:=\operatorname{Tr}(\ket{\psi(x,\theta)}\bra{\psi(x,\theta)}\mathcal{M}_{\theta}),$
$\mathcal{H}$ :=$:=\{h(x,\theta):\theta\in\Theta\},$
with $\Theta$ being the parameter space. Typically, $U(x,\theta)$ incorporates data $x$ and parameters $\theta$ through rotational gates as in Figure 2, leading to Fourier series models in certain instances [^1] This is the case if we do not use feature-dependent multi-qubit gates for the data embedding, see also Schuld et al. (2021).:
$h(x,\theta)=\sum_{\omega}c_{\theta}^{\omega}e^{-\mathrm{i}\omega\cdot x}.$ (2)
Using PPTNQFE, we obtain the following expression for the model
$h(x,\theta)$ $=\braket{\Phi(x)}{U^{\dagger}(\theta)\mathcal{M}_{\theta}U(\theta)}{\Phi(x)}$
$=(n_{\Phi}(x))^{-2}\sum_{\bar{k},k}\varphi_{\bar{k}}(x)\varphi_{k}(x)\braket{\bar{k}}{U^{\dagger}(\theta)\mathcal{M}_{\theta}U(\theta)}{k}$
$=2(n_{\Phi}(x))^{-2}\sum_{\bar{k}\geq k}\varphi_{\bar{k}}(x)\varphi_{k}(x)\Re\braket{\bar{k}}{U^{\dagger}(\theta)\mathcal{M}_{\theta}U(\theta)}{k}.$ (3)
Note that for the $x$-dependence we obtain products $\varphi_{\bar{k}}(x)\varphi_{k}(x)$. To include the original piecewise polynomials from Section 2 within this functional space, we set $\varphi_{k}:=\sqrt{\hat{\varphi}_{k}}$. Due to property (III) this automatically normalizes the basis vector so that $n_{\Phi}(x)=1$. The entire set of functions $\varphi_{\bar{k}}\varphi_{k}$ is plotted in Figure 4.
Figure 4: Plots of products of basis functions $\varphi_{\bar{k}}\varphi_{k}$, $\bar{k}\geq k$, on 2 qubits (10 products in total).
It is crucial to recognize that these data embeddings are not inherently superior or inferior to conventional rotation-based embeddings; they merely offer an alternative functional space that may better suit specific applications. This diversity in embeddings introduces different inductive biases into the models.
To illustrate this point, consider Figure 5. These are plots of $h(x,\theta=0)$ for $x\in[-1,1]$ and different observables $\mathcal{M}$, i.e., the data embedding contains no trainable layers and only parameter-independent observables. From (5.2), we see that observables diagonal in the computational $Z$-basis correspond to combinations of piecewise polynomials $\hat{\varphi}_{k}$ only. The qubit number on which we measure reflects the frequency of sign changes. Vice versa, purely off-diagonal observables, like the Pauli-$X$, correspond to non-polynomial elements $\sqrt{\hat{\varphi}_{\bar{k}}\hat{\varphi}_{k}}$, $\bar{k}\neq k$. Similar statements hold for products of these observables.
Figure 5: Plots of only PPTNQFE and different observables, no variational layers or trainable parameters.
In Figure 6, we approximate a discontinuous jump function with added Gaussian noise. Traditional rotation-based embeddings will struggle with such functions, necessitating high frequencies to capture discontinuities. On the contrary, PPTNQFE can achieve a satisfactory approximation with two trainable parameters. For PPTNQFE, we use no variational layers and a parameter-dependent observable $\mathcal{M}_{\theta}=\theta_{0}I+\theta_{1}Z_{0}$. For the rotation-based model, we use exponential feature encoding as in Figure 2, the same observable and 3 variational layers, resulting in a total of 89 trainable parameters.
Figure 6: The training samples are based on the function top left. The PPTNQFE approximation is top right. The rotation-based model is on the bottom.
Next, consider a $\sin$ wave perturbed by Gaussian noise. For such a function, a rotation-based model with a limited frequency spectrum would generally provide an adequate approximation. However, to demonstrate overfitting, we employ exponential encoding in the rotation-based embedding, as illustrated in Figure 2. We attempt fitting the rotation-based model on 5 qubits with a
  • a single variational layer consisting of universal single-qubit gates and CNOTs, combined with 5 trainable observables $\theta_{k}Z_{k}$, and a constant bias term, culminating in a total of 45 trainable parameters;
  • and an expanded setup with 3 variational layers, increasing the total count of trainable parameters to 93.
For PPTNQFE, we use an MPS-structured variational layer, similar to Figure 2, but with three 2-qubit gates instead of a single layer. The rationale for employing an MPS-structured variational layer is based on the idea that the function $\sin$ can be represented as a rank-2 MPS, as detailed in Oseledets (2013). By aligning the variational layer with a rank-2 MPS structure, the aim is to discover a unitary transformation that effectively maps back to the computational basis. This method has shown improved outcomes compared to the use of generic variational layers, although it remains a heuristic approach at this stage. We use observables similar to those in the rotation-based model, $\mathcal{M}_{\theta}=\theta_{0}+\sum_{k}\theta_{k+1}Z_{k}$, resulting in 93 trainable parameters in total. The outcomes are presented in Figure 7.
Figure 7: The training samples are based on the function top left. The PPTNQFE approximation with an MPS variational layer (93 parameters) is top right. The rotation-based model with 3 variational layers (93 parameters) is bottom left. The rotation-based model with 1 variational layers (45 parameters) is bottom right.
Both rotation-based models overfit noise, with the simpler one performing slightly better. This is mainly due to the accessible spectrum being too large and no bias towards the correct solution. The number of parameters was deliberately taken to be larger than necessary to demonstrate noise overfitting. Similarly, PPTNQFE with 93 parameters does not perform as well as for the jump function, with square root cusps introduced due to the added variational layers in the model.
Given a budget of $n$ qubits and using PPTNQFE, we can encode $2^{n}$ piecewise polynomials and, taking into account off-diagonal terms, $2^{2n-1}+2^{n-1}$ non-polynomial terms. Measuring simple observables such as $Z_{0}$ means, the final approximation is a linear combination of all piecewise polynomials. If we want a local approximation, addressing each piecewise polynomial $\hat{\varphi}_{k}$ individually, doing so naively would require measuring
$\mathcal{M}_{\theta}:=\sum_{k}\theta_{k}\ket{k}\bra{k},$
with $2^{n}$ trainable parameters. This is similar to the rotation-based models from (2), where a limited number of parameters means we only have access to a small subset of the Fourier coefficients $\{c_{\theta}^{\omega}\}_{\omega}$. There is no “free lunch” in either case, and, despite the ambient Hilbert space being exponentially large, a good hypothesis class $\mathcal{H}$ for the given learning problem will require choosing the right structure for the trainable unitaries and observables.
As an example, consider the observable $\mathcal{M}_{\theta}:=\theta Z_{0}$, and the target function from Figure 8 with 2 discontinuous jumps. Clearly, a sufficiently good approximation would only take the first two and the last two piecewise polynomials $\hat{\varphi}_{k}$, $k=0,1,2^{n}-2,2^{n}-1$. From (5.2), we see that this would require a unitary that rotates into a basis in which the observable $Z_{0}$ has diagonal entries $1,1,0,\ldots,0,-1,-1$ and arbitrary, but purely complex-valued, off-diagonal entries. As can be seen in Figure 8, this is indeed achievable. Here, the unitary was obtained analytically by direct optimization, see Section 7. More generally, problem-specific, efficient quantum circuit designs is an interesting open question.
Figure 8: The training samples are based on the function left. The PPTNQFE model appends a non-trainable layer based on an analytically computed unitary and measures $\mathcal{M}_{\theta}=\theta Z_{0}$.
6 Conclusions
In this work, we introduced a methodology to embed continuous variables in quantum circuits using piecewise polynomial features. This approach involves representing a piecewise polynomial basis as a low-rank TN, which is then transformed into a quantum circuit. The conversion process results in a quantum circuit where the number of qubits for the converted TN quantum gates scales logarithmically with the rank of the TN, making the embeddings potentially efficient for low-rank TNs with the number of gates growing linearly with the qubit count.
However, converting TN gates to those native to quantum hardware may introduce significant overhead. Additionally, the effects of statistical and hardware-induced noise on these embeddings were not explored in this work.
We explored two applications of this method in one-dimensional settings. First, PPTNQFE allows efficient point evaluations of discretized solutions to PDEs, enhancing the appeal of quantum algorithms for solving PDEs by addressing data read-out challenges. Second, we demonstrated the utility of PPTNQFE in function regression, particularly highlighting its aptitude for modeling functions with localized features, such as jump discontinuities. An interesting open question is how to design problem-specific trainable circuits suited for approximating more complex dependencies with local effects.
The broader objective of our research is to widen the utility of quantum models by introducing a novel approach to feature embeddings that takes advantage of TN representations. This not only facilitates the embedding of piecewise polynomials, but also opens avenues for incorporating various other function types into quantum models. The rich existing literature on the theoretical and numerical aspects of TN representations could provide valuable insights, contributing to a more profound understanding of the potential of QML. We aim to unveil the inner mechanics of quantum models in a transparent manner, focusing on comprehending how design choices impact their functionality, the scope of problems they are suited for, and exploring their potential benefits, if any.
7 Data Availability
The source code for all the numerical experiments presented here can be found at https://github.com/MazenAli/PPTNQFE_Examples. Running the notebooks requires installing the package https://github.com/MazenAli/QuLearn.
8 Acknowledgments
We thank BMWK for the financial support provided under the EniQmA project and the Competence Center for Quantum Computing Rhineland-Palatinate for the financial support provided under the AnQuC-3 project.

### v5

@@
 This work introduces a novel method for embedding continuous variables into quantum circuits via piecewise polynomial features, utilizing low-rank tensor networks.
-Our approach, termed Piecewise Polynomial Tensor Network Quantum Feature Encoding (PPTNQFE), aims to broaden the applicability of quantum algorithms by incorporating spatially localized representations suited for numerical applications like solving partial differential equations and function regression.
+Our approach, termed Piecewise Polynomial Tensor Network Quantum Feature Encoding (PPTNQFE), aims to broaden the applicability of quantum algorithms by incorporating spatially localized representations suited for numerical applications like partial differential equations and function regression.
 We demonstrate the potential of PPTNQFE through efficient point evaluations of solutions of discretized differential equations and in modeling functions with localized features such as jump discontinuities.
-Although promising, challenges such as unexplored noise impact and design of trainable circuits remain.
-This study opens new avenues for enhancing quantum models with novel feature embeddings and leveraging TN representations for a wider array of function types in quantum machine learning.
+While promising, challenges such as unexplored noise impact and design of trainable circuits remain.
 Keywords Quantum Machine Learning  $\cdot$ Feature Encoding  $\cdot$ Tensor Networks  $\cdot$ Piecewise Polynomial  $\cdot$ Splines  $\cdot$ Differential Equations  $\cdot$ Finite Element Methods
@@
 The rationale for employing PQCs often hinges on the large size of the quantum state space.
-However, these circuits often produce models similar to Fourier series, which may not fully take advantage of the exponentially large Hilbert space and can be emulated classically Schuld et al.
+Nonetheless, these circuits often produce Fourier series models, which may not fully take advantage of the exponentially large Hilbert space and can be emulated classically Schuld et al.
 (2021); Kordzanganeh et al.
@@
 The efficacy of a model extends beyond mere expressiveness; it depends on the synergy between sufficient expressiveness and the appropriate, problem-specific inductive bias.
-Recent innovative studies have explored the potential to imbue models with biases that are distinctively quantum in nature Bowles et al.
+Recent innovative studies have explored the potential to encode specific biases more efficiently using quantum models Bowles et al.
 (2023); Gili et al.
@@
 (2023).
-Our approach leverages an intermediate tensor network (TN) representation, which makes it adaptable to any function that can be represented in low-rank format Oseledets (2013); Ali and Nouy (2023).
+Our approach leverages an intermediate tensor network (TN) representation, which makes it adaptable to any function that can be represented in a low-rank format Oseledets (2013); Ali and Nouy (2023).
 We denote these embeddings as Piecewise Polynomial Tensor Network Quantum Feature Encodings (PPTNQFE).
@@
 Section 3 discusses TNs and functional TN representations.
-Section 4 details the methodology for expressing piecewise polynomial bases as low-rank TNs and their subsequent translation into quantum circuits.
-Section 5 presents two important applications: point evaluations of PDE solutions and function learning (regression).
+Section 4 details the methodology for expressing piecewise polynomial bases as low-rank TNs and their subsequent conversion into quantum circuits.
+Section 5 presents three important applications: point evaluations of PDE solutions, function learning (regression) and classification with quantum kernels.
 The paper concludes in Section 6.
-The implementation and examples provided are confined to one spatial dimension, with further details in Section 7.
-While extending these methods efficiently to higher dimensions is theoretically feasible, it would entail a more involved implementation.
+All computations with quantum algorithms were performed using a state vector simulator.
 2 Piecewise Polynomials
@@
   • In Section 5.2, $x_{0}=-1$, $x_{2^{n}-1}=1$.
-A plot of the hat functions is presented in Figure 1. These satisfy conditions (I), (II) and (III).
-Figure 1: 1D piecewise linear hat basis for a grid of 8 uniformly spaced points, including the boundary $x\in\{\pm 1\}$
+A plot of the hat functions is presented in Figure 1.
+These satisfy conditions (I), (II) and (III).
+Moreover, by taking tensor products of the proposed feature embeddings, we can trivially extend them to any spatial dimension.
+See Figure 2 for a two-dimensional example.
+Figure 1: 1D piecewise linear hat basis for a grid of 8 uniformly spaced points, including the boundary $x\in\{\pm 1\}$.
+Figure 2: 2D bilinear basis functions formed as tensor products of 4 linear basis functions in each coordinate direction, i.e., 16 functions in total.
 3 Tensor Networks
@@
 In this work, we will be using the matrix product state (MPS) format Schollwöck (2011), also known under the name tensor train (TT) format Oseledets (2011), see Figure 3.
+A state |$\ket{\psi}$ over $n$ sites is called an MPS (with open boundary conditions) if it can be expressed as
+|$\ket{\psi}=\sum_{s_{0},\ldots,s_{n-1}}C^{0}_{s_{0}}\cdots C^{n-1}_{s_{n-1}}\ket{s_{0}\ldots s_{n-1}}.$
 Figure 3: On the top left, an example of a rotation-based exponential encoding.
@@
 $C^{n-1}(x)[1,1,0]$ $=a.$
-The case of even $k$ is similar. Since the maximum rank is $2$, converting to a quantum circuit as in Figure 2. we require only $2$-qubit gates.
+The case of even $k$ is similar. Since the maximum rank is $2$, converting to a quantum circuit (see Figure 3) we require only $2$-qubit gates.
 Remark 4.1
@@
 (2024): here we load the set of functions $\{\varphi_{k}(x)\}_{k=0}^{N-1}$, evaluated at a point $x\in\overline{\Omega}$, instead of loading a discrete representation of a single function $\{f_{k}=f(x_{k})\}_{k=0}^{N-1}$.
+Finally, having embedded 1D features, we can easily extend this approach to any number of features by taking tensor products of the MPSs.
+For instance, given three features $x$, $y$, and $z$ with corresponding binary indices $i_{k}$, $j_{k}$, and $m_{k}$, one possible construction is the canonical ordering:
+$\Phi(x,y,z)=$ $A^{0}(x)[\cdot,i_{0},\cdot]\cdots A^{n-1}(x)[\cdot,i_{n-1},\cdot]\cdot$
+$B^{0}(y)[\cdot,j_{0},\cdot]\cdots B^{n-1}(y)[\cdot,j_{n-1},\cdot]\cdot$
+$C^{0}(z)[\cdot,m_{0},\cdot]\cdots C^{n-1}(z)[\cdot,m_{n-1},\cdot].$
+Another commonly used alternative is the transposed ordering
+$\Phi(x,y,z)=$ $A^{0}(x)[\cdot,i_{0},\cdot]\cdot B^{0}(y)[\cdot,j_{0},\cdot]\cdot C^{0}(z)[\cdot,m_{0},\cdot]\cdots$
+$A^{n-1}(x)[\cdot,i_{n-1},\cdot]\cdot B^{n-1}(y)[\cdot,j_{n-1},\cdot]\cdot C^{n-1}(z)[\cdot,m_{n-1},\cdot].$
+For further details on the effects of different orderings, see Kazeev et al.
+(2017); Markeeva et al.
+(2021); Ali and Nouy (2023).
+All orderings yield a feature vector $\Phi(x,y,z)$ with the same values but arranged in different sequences.
+Remark 4.2
+We emphasize that constructing the MPS for this feature encoding is computationally efficient, with a complexity of $\mathcal{O}(r)$.
+Specifically, in the example discussed in this paper, where $r=2$, the construction is particularly inexpensive.
 5 Applications
-In this section, we dive into two illustrative applications of the PPTNQFE method that we have introduced.
+In this section, we dive into three illustrative applications of the PPTNQFE method that we have introduced.
 The first application focuses on the efficient point evaluation of solutions to differential equations.
@@
 Here, PPTNQFE allows, e.g., the approximation of jump functions in a particularly simple manner.
+Finally, the third application is classification with quantum kernels.
+We demonstrate on synthetic datasets that, in some cases, PPTNQFE achieves similar results to rotation-based encoding.
+In other cases, PPTNQFE can avoid misclassification due to the periodic nature of rotation-based encodings.
 5.1 Differential Equations
@@
 $u(x)$ $=0,$ $\quad x\in\{\pm 1\}.$
-The associated linear system, represented in matrix form,
-$\boldsymbol{A}\boldsymbol{u}=\frac{1}{h}\begin{pmatrix}2&-1&0&\dots&\\ -1&2&-1&\dots&\\ &&\ddots&&\\ &&&&-1\\ &&&-1&2\end{pmatrix}\boldsymbol{u}=\boldsymbol{f}.$
+The associated (rescaled) linear system, represented in matrix form,
+$\boldsymbol{A}\boldsymbol{u}=\begin{pmatrix}2&-1&0&\dots&\\ -1&2&-1&\dots&\\ &&\ddots&&\\ &&&&-1\\ &&&-1&2\end{pmatrix}\boldsymbol{u}=\boldsymbol{f}.$
 can be solved using a variational quantum algorithm.
@@
 In the context of supervised learning, we are presented with a training set $\mathcal{S}_{M}:=\left\{(x_{0},y_{0}),\ldots,(x_{M},y_{M})\right\}\in(X\times Y)^{M}$ consisting of independently and identically distributed samples from an unknown distribution $\mathcal{P}(X\times Y)$ and a hypothesis class $\mathcal{H}$.
-The objective is to identify a model $h\in\mathcal{H}$, $h:X\rightarrow Y$, that minimizes the empirical risk
-$R_{M}(h):=\frac{1}{M}\sum_{i}L(h(x_{i}),y_{i}),$
+The objective is to identify a model $h\in\mathcal{H}$, $h:X\times\Theta\rightarrow Y$, that minimizes the empirical risk
+$R_{M}(h):=\frac{1}{M}\sum_{i}L(h(x_{i},\theta),y_{i}),$
 where $L:Y\times Y\rightarrow\mathbb{R}$ is a specified loss function.
@@
 $=(n_{\Phi}(x))^{-2}\sum_{\bar{k},k}\varphi_{\bar{k}}(x)\varphi_{k}(x)\braket{\bar{k}}{U^{\dagger}(\theta)\mathcal{M}_{\theta}U(\theta)}{k}$
-$=2(n_{\Phi}(x))^{-2}\sum_{\bar{k}\geq k}\varphi_{\bar{k}}(x)\varphi_{k}(x)\Re\braket{\bar{k}}{U^{\dagger}(\theta)\mathcal{M}_{\theta}U(\theta)}{k}.$ (3)
+$=2(n_{\Phi}(x))^{-2}\sum_{\bar{k}\geq k}\varphi_{\bar{k}}(x)\varphi_{k}(x)\Re\braket{\bar{k}}{U^{\dagger}(\theta)\mathcal{M}_{\theta}U(\theta)}{k},$ (3)
+for any unitary $U(\theta)$.
 Note that for the $x$-dependence we obtain products $\varphi_{\bar{k}}(x)\varphi_{k}(x)$.
@@
 These are plots of $h(x,\theta=0)$ for $x\in[-1,1]$ and different observables $\mathcal{M}$, i.e., the data embedding contains no trainable layers and only parameter-independent observables.
-From (5.2), we see that observables diagonal in the computational $Z$-basis correspond to combinations of piecewise polynomials $\hat{\varphi}_{k}$ only.
+From (5.2), we see that expectations values of operators that are diagonal in the computational $Z$-basis correspond to combinations of piecewise polynomials $\hat{\varphi}_{k}$ only.
 The qubit number on which we measure reflects the frequency of sign changes.
@@
 Similar statements hold for products of these observables.
-Figure 5: Plots of only PPTNQFE and different observables, no variational layers or trainable parameters.
+Figure 6: Plots of expectation values of different observables, applied immediately after the PPTNQFE embedding – no variational layers or trainable parameters.
+The following training of quantum regression models was performed on quantum state vector simulator with AMSGrad, using the MSE loss on a training set of 50 equally spaced points.
+More details can be found in the notebooks, see Section 7.
 In Figure 7, we approximate a discontinuous jump function with added Gaussian noise.
@@
 On the contrary, PPTNQFE can achieve a satisfactory approximation with two trainable parameters.
-For PPTNQFE, we use no variational layers and a parameter-dependent observable $\mathcal{M}_{\theta}=\theta_{0}I+\theta_{1}Z_{0}$.
+For PPTNQFE, we use no variational layers and a parameter-dependent observable $\mathcal{M}_{\theta}=\theta_{0}I+\theta_{1}Z_{0}$, where $Z_{0}$ is the $Z$-operator applied to the 00’th qubit.
 For the rotation-based model, we use exponential feature encoding as in Figure 3, the same observable and 3 variational layers, resulting in a total of 89 trainable parameters.
@@
 We attempt fitting the rotation-based model on 5 qubits with a
-  • a single variational layer consisting of universal single-qubit gates and CNOTs, combined with 5 trainable observables $\theta_{k}Z_{k}$, and a constant bias term, culminating in a total of 45 trainable parameters;
+  • a single variational layer consisting of universal single-qubit gates and CNOTs, combined with 5 trainable observables $\theta_{k}Z_{k}$ ($Z_{k}$ is the $Z$-operator applied to qubit $k$), and a constant bias term, culminating in a total of 45 trainable parameters;
   • and an expanded setup with 3 variational layers, increasing the total count of trainable parameters to 93.
 For PPTNQFE, we use an MPS-structured variational layer, similar to Figure 3, but with three 2-qubit gates instead of a single layer.
-The rationale for employing an MPS-structured variational layer is based on the idea that the function $\sin$ can be represented as a rank-2 MPS, as detailed in Oseledets (2013).
-By aligning the variational layer with a rank-2 MPS structure, the aim is to discover a unitary transformation that effectively maps back to the computational basis.
+The rationale for employing an MPS-structured variational layer is based on the idea that the $\sin$ function can be represented as a rank-2 MPS, as detailed in Oseledets (2013).
+By using a variational layer with a similar structure to the PPTNQFE embedding layer, the aim is to discover a unitary transformation that effectively maps back to the computational basis.
 This method has shown improved outcomes compared to the use of generic variational layers, although it remains a heuristic approach at this stage.
@@
 Similarly, PPTNQFE with 93 parameters does not perform as well as for the jump function, with square root cusps introduced due to the added variational layers in the model.
+The occurrence of non-polynomial terms was explained in Figure 6.
 Given a budget of $n$ qubits and using PPTNQFE, we can encode $2^{n}$ piecewise polynomials and, taking into account off-diagonal terms, $2^{2n-1}+2^{n-1}$ non-polynomial terms.
@@
 The PPTNQFE model appends a non-trainable layer based on an analytically computed unitary and measures $\mathcal{M}_{\theta}=\theta Z_{0}$.
+5.3 Classification
+For the final example, we consider classification with quantum kernels.
+As pointed out in Schuld and Petruccione (2021), supervised QML methods without data re-uploading are kernel methods.
+Thus, we can determine the optimal solution by minimizing the hinge loss over a training set through inverting the kernel matrix.
+If $x\mapsto\rho(x)$ is the embedding map, the kernel map is given by
+$\kappa(x,x^{\prime})=\operatorname{Tr}[\rho(x^{\prime})\rho(x)],$
+and the prediction for any new point $x^{*}$ is computed as
+$f(x^{*})=\sum_{m=1}^{M}\alpha_{m}\kappa(x^{*},x_{m}),$
+where $M$ is the number of training samples and the coefficients $\alpha_{m}$ are determined during training.
+We apply this method to both the angle-based embedding and PPTNQFE, and test the quantum kernel classification on two synthetic datasets, as shown in Figures 10 and 11.
+For a fair comparison, we use 6 qubits for both models.
+For the angle-based embedding, this corresponds to a Fourier spectrum of 13 different frequencies per feature (169 in total), and for PPTNQFE, this corresponds to 8 hat functions for each feature (64 in total).
+Both models perform similarly well on the two half moons dataset. Some points are misclassified due to noise, which places some points of different classes very close to each other.
+For the line-like clusters in Figure 11, a simple linear model should already perform well.
+However, the angle-based embedding misclassifies some points that are far from the clusters due to the periodic nature of the embedding and the presence of high-frequency components in the model.
+On the other hand, PPTNQFE does not suffer from this issue and only misclassifies one point that is close to the boundary between the two clusters.
+Figure 10: Top left: full dataset of two half moons representing two classes with noise.
+The training set for both quantum kernels was chosen as 80% of the full dataset, randomly selected.
+Top right: classification results on the full dataset for the angle-based embedding after training.
+Bottom: classification results on the full dataset for PPTNQFE after training.
+Figure 11: Top left: full dataset of three lines representing two classes with noise.
+The training set for both quantum kernels was chosen as 80% of the full dataset, randomly selected.
+Top right: classification results on the full dataset for the angle-based embedding after training.
+Bottom: classification results on the full dataset for PPTNQFE after training.
 6 Conclusions
@@
 Additionally, the effects of statistical and hardware-induced noise on these embeddings were not explored in this work.
-We explored two applications of this method in one-dimensional settings.
+We explored three applications of this method.
 First, PPTNQFE allows efficient point evaluations of discretized solutions to PDEs, enhancing the appeal of quantum algorithms for solving PDEs by addressing data read-out challenges.
 Second, we demonstrated the utility of PPTNQFE in function regression, particularly highlighting its aptitude for modeling functions with localized features, such as jump discontinuities.
+Third, we showed that PPTNQFE can easily identify non-linear decision boundaries, similar to angle-based embeddings, while avoiding some of the typical issues associated with the periodic Fourier nature of angle-based embeddings.
 An interesting open question is how to design problem-specific trainable circuits suited for approximating more complex dependencies with local effects.